# Interval Decomposition of Infinite Persistence Modules over a Principal Ideal Domain

### v1

@@ 1. Introduction
Let $R$ be a principal ideal domain (PID). A persistence module with coefficients in $R$ is a functor $f:I\to R\mathrm{-Mod}$, where the indexing category $I$ is a totally-ordered poset category. Persistence modules arise from and have been widely studied in persistent homology, in which setting $R$ is typically a field and $f$ maps every index $i\in I$ to a finite-dimensional vector space over $R$. In this setting, any persistence module admits an interval decomposition; that is, it can be decomposed as a direct sum of indecomposables known as “interval modules”. Interval decompositions are key to persistent homology, in that they capture important topological structure within data. If one removes the condition that the coefficient ring $R$ be a field, then interval decompositions are no longer guaranteed to exist.
@@
The existence of interval decompositions for modules with field coefficients has been widely attributed to Gabriel [6], whose work showed that indecomposable representations of quivers of type $A_{n}$ have only finitely many isomorphism classes, which coincide exactly with interval modules. This result has since been generalized to persistence modules indexed by any totally-ordered set [4], as well as a small category [2]. Interval decompositions have also been generalized to zigzag persistence. Carlson and de Silva [3] showed that interval decompositions exist for finitely-indexed zigzag persistence modules; Botnan [1] later showed the same in the infinitely-indexed setting. Hang and Mio [9] generalized interval decomposition to tame correspondence modules.
@@
Our work complements and builds on [10], which established necessary and sufficient conditions for interval decompositions of persistence modules over PIDs with finite indexing categories. The core element of [10] is a polynomial-time algorithm to compute decompositions explicitly. While key ideas in [10] remain relevant, the generality of our indexing category requires nontrivial adaptation.
@@ 2.1. Persistence modules
We first show that the family $\{\mathrm{Ker}[a,y]\}_{y\in I}$ of kernels out of $f_{a}$ is finite. Recall that for $y\leq a$, we define $\mathrm{Ker}[a,y]=0$, while for $y>a$ we define $\mathrm{Ker}[a,y]=\ker(f(a\leq y))$. Let $y_{1},y_{2}\in I$ such that $a<y_{1}\leq y_{2}$. Then we have a short exact sequence
@@
We now show that the family $\{\mathrm{Im}[x,a]\}_{x\in I}$ of images into $f_{a}$ is finite. Recall that for $x\geq a$, we define $\mathrm{Im}[x,a]=f_{a}$, while for $x<a$ we define $\mathrm{Im}[x,a]=\mathrm{Im}(f(x\leq a))$. Let $x_{1},x_{2}\in I$ be such that $x_{1}\leq x_{2}<a$. Then we have a short exact sequence
@@ 2.2. A brief discussion on finitely-indexed persistence modules
If the structure maps of $f$ all have free cokernels, then [10, Theorem 17] states that, for each pair $(p,q)$ of integers such that $1\leq p,q\leq n$, there exits a (non-unique) submodule $A^{pq}_{a}\subseteq f_{a}$ such that
@@
Theorem [10, Theorem 18] states that, independent of one’s choice of $A^{pq}_{a}$, one has
@@
In the context of persistent homology, a submodule $A_{a}^{pq}\subseteq f_{a}$ in this decomposition can be thought of as cycles that are born at $p$ and die at $q$. This is because (1) whenever $a\notin[p,q)$, we must have $A_{a}^{pq}=0$ (see [10, Lemma 15]), and (2) one can choose submodules $A^{pq}_{a}$ such that $f(a\leq b)(A_{a}^{pq})=A_{b}^{pq}$. That is, for any choice of $A_{a}^{pq}$, the image $f(a\leq b)(A_{a}^{pq})$ is a complement of $\mathrm{IK}_{b}[p-1,q]+\mathrm{IK}_{b}[p,q-1]$ in $\mathrm{IK}_{b}[p,q]$ (see [10, Lemma 16]).
@@ 4. Sufficiency
Throughout this section, we fix a PID $R$, our coefficent ring; fix a persistence module $f:I\to R\mathrm{-Mod}$ that is pointwise free and finitely-generated; assume that every structure map has free cokernel.
@@ 4.2. Partitioning consistent bases with images and kernels
Fix $J\subsetneq I$ and $a\in I\setminus J$. Choose a finite index set $S=\{i_{1},\ldots,i_{r}\}\subseteq I$ such that, for every $c\in I$, we have $\mathrm{Ker}[a,c]=\mathrm{Ker}[a,i_{t}]$ for some $t$ and $\mathrm{Im}[c,a]=\mathrm{Im}[i_{s},a]$ for some $s$. Such a finite set exist by Lemma 2.9. We further require $\{i_{1},\ldots,i_{r}\}$ to (1) contain $a$, (2) be of minimal size, and (3) contain as many elements of $J$ as possible. (That is, for any $j\in J$ such that $j\geq a$, there must be some $i_{q}\in J$ such that $\mathrm{Ker}[a,j]=\mathrm{Ker}[a,i_{q}]$. Similarly, for any $j\in J$ such that $j\leq a$, there must be some $i_{p}\in J$ such that $\mathrm{Im}[j,a]=\mathrm{Im}[i_{p},a]$.)
@@ 4.3. Incremental Extension
    If none of the preceding rules applies (i.e., $i_{s}<i_{p}\leq a<i_{q}\leq i_{t}$), we choose any complement $A_{a}^{pq}$ of $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}[i_{p},i_{q-1}]$ in $\mathrm{IK}_{a}[i_{p},i_{q}]$, and define $\beta^{pq}_{a}$ to be any basis of $A_{a}^{pq}$. (Such a complement exists by [10, Theorem 17].) In this case, there is nothing in $J$ to push forward or pull back to index $a$, so we choose an arbitrary complement.
@@
In case (2), we have, by [10, Lemma 16], that $f(i_{s}\leq a)|_{{A_{i_{s}}^{pq}}}$ restricts to an isomorphism onto its image. Because $\beta^{pq}_{i_{s}}$ is linearly independent (it is a basis of $A^{pq}_{i_{s}}$), its image $f(i_{s}\leq a)(\beta^{pq}_{i_{s}})$ is as well.
@@
Case (1) Suppose that $i_{p}>a$ or $i_{q}\leq a$. By the proof of [10, Lemma 15], we have that $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]=\mathrm{IK}_{a}[i_{p},i_{q}]$. This implies that the complement is 0, with basis $\beta^{pq}_{a}:=\emptyset$, as desired.
Case (2) Suppose that $p\leq s$. By [10, Lemma 16], $f(i_{s}\leq a)(A_{i_{s}}^{pq})$ is a complement of $\mathrm{IK}_{a}[i_{p},i_{q-1}]+\mathrm{IK}_{a}[i_{p-1},i_{q}]$ in $\mathrm{IK}_{a}[i_{p},i_{q}]$.
@@
Figure 1. The four cases for defining $\beta_{a}^{pq}$, for different values of $i_{p}$ and $i_{q}$.
@@
To show that $\beta_{a}=\bigsqcup_{p,q}\beta^{pq}_{a}$ is a basis for $f_{a}$, it is enough to show that $f_{a}=\bigoplus_{p,q}A^{pq}_{a}$. This is given by [10, Theorem 18]. ∎
@@
    $i_{p}>i_{s}$ or $i_{q}\leq i_{s}$. In this case, the basis $\beta^{pq}_{i_{s}}$ is empty, by [10, Lemma 15].
@@
Corollary 4.12 (Matching matrices for $(S\cap J)\cup\{a\}$).
@@
The matrix representation of every structure map with respect to the bases $(\beta_{j})_{j\in J\cup\{a\}}$ is a matching matrix. That is, for every $c,b\in J\cup\{a\}$, the matrix representation $f(c\leq b)]_{\beta_{c}}^{\beta_{b}}$ is a matching matrix.
@@
We first look at maps that end at $f_{a}$. Take any $c\in J$ such that $c<a$. Choose $i_{u}\in S\cap J$ such that $\mathrm{Im}[i_{u},a]=\mathrm{Im}[c,a]$. If $c\leq i_{u}\leq a$, then $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}=[f(i_{u}\leq a)]_{\beta_{i_{u}}}^{\beta_{a}}[f(c\leq i_{u})]_{\beta_{c}}^{\beta_{i_{u}}}$. Note that $f(c\leq i_{u})]_{\beta_{c}}^{\beta_{i_{u}}}$ is a matching matrix because $i_{u},c\in J$, and $(\beta_{j})_{j\in J}$ is a consistent basis. Additionally, by Corollary 4.12, $f(i_{u}\leq a)]_{\beta_{i_{u}}}^{\beta_{a}}$ is a matching matrix. Therefore $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix.
Now, consider $i_{u}\leq c<a$, and let $\{v_{1},\ldots,v_{r}\}:=\beta_{i_{u}}\setminus\mathrm{Ker}[i_{u},a]$. By Corollary 4.12, this set maps injectively into $\beta_{a}$, and its image under $f(i_{u}\leq a)$ spans $\mathrm{Im}[i_{u},a]=\mathrm{Im}[c,a]$. By functoriality, this implies that the image of the set $\{f(i_{u}\leq c)(v_{1}),\ldots,f(i_{u}\leq c)(v_{r})\}\subseteq\beta_{c}$ maps injectively into the basis $\beta_{a}\subseteq f_{a}$ under $f(c\leq a)$. Because $\beta_{c}$ contains a basis for $\mathrm{Ker}[c,a]$, we have, by the rank-nullity theorem, that $\beta_{c}\setminus\{f(i_{u}\leq c)(v_{1}),\ldots,f(i_{u}\leq c)(v_{r})\}$ spans $\mathrm{Ker}[c,a]$. Therefore, it follows that $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix.
Next, we look at maps that begin at $f_{a}$. Take any $b\in J$ such that $b>a$. By definition, there exists $i_{\ell}\in S\cap J$ such that $\mathrm{Ker}[a,i_{\ell}]=\mathrm{Ker}[a,b]$. If $a<i_{\ell}\leq b$, then $f(a\leq b)]_{\beta_{a}}^{\beta_{b}}=[f(i_{\ell}\leq b)]_{\beta_{i_{\ell}}}^{\beta_{b}}[f(a\leq i_{\ell})]_{\beta_{a}}^{\beta_{i_{\ell}}}$ is a matching matrix because both $f(i_{\ell}\leq b)]_{\beta_{i_{\ell}}}^{\beta_{b}}$ and $f(a\leq i_{\ell})]_{\beta_{a}}^{\beta_{i_{\ell}}}$ (by Corollary 4.12) are matching matrices.
Now, suppose that $a<b\leq i_{\ell}$. Let $v\in\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]$ be given. Because $\mathrm{Ker}[a,i_{\ell}]=\mathrm{Ker}[a,b]$, the elements $f(a\leq b)(v)$ and $f(a\leq i_{\ell})(v)$} are nonzero. We know that $f(a\leq i_{\ell})(v)\in\beta_{i_{\ell}}$, by Corollary 4.12. Therefore, because $(\beta_{j})_{j\in J}$ is a consistent basis, there exists $w\in\beta_{b}$ such that $f(b\leq i_{\ell})(w)=f(a\leq i_{\ell})(v)$.
@@
From this, it follows that $f(a\leq b)$ maps $\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]=\beta_{a}\setminus\mathrm{Ker}[a,b]$ into $\beta_{b}$. Moreover, functoriality implies that this map is injective, because the composite map $f(a\leq i_{\ell})=f(b\leq i_{\ell})\circ f(a\leq b)$ is injective on $\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]=\beta_{a}\setminus\mathrm{Ker}[a,b]$, by Corollary 4.12. This implies that $f(a\leq b)]_{\beta_{a}}^{\beta_{b}}$ is a matching matrix, as desired. ∎
@@
Suppose that $g:K\to R\mathrm{-Mod}$ is a persistence module with consistent basis $(\gamma_{k})_{k\in K}$. For any $a\in K$, the basis $\gamma_{a}$ contains a basis for every submodule of the form $\mathrm{Im}[c,a]$ and $\mathrm{Ker}[a,b]$.
@@
For any $x,y\in K$ such that $x\leq y$, define $\mathcal{K}_{x,y}=\gamma_{x}\cap\mathrm{Ker}[x,y]$, and $\mathcal{I}_{x,y}=\gamma_{x}\setminus\mathcal{K}_{x,y}$. Let $c\leq a\leq b$ be given. Clearly, $f(c\leq a)(\mathcal{I}_{c,a})$ spans the image submodule $\mathrm{Im}[c,a]$. Because $f(c\leq a]_{\gamma_{c}}^{\gamma_{a}}$ is a matching matrix, the set $\mathcal{I}_{c,a}$ maps injectively into $\gamma_{a}$, and therefore, $f(c\leq a)(\mathcal{I}_{c,a})\subseteq\gamma_{a}$ forms a basis for $\mathrm{Im}[c,a]$. Moreover, by the rank nullity theorem, $\mathcal{N}_{a,b}$ forms a basis for $\mathrm{Ker}[a,b]$. ∎
@@
Corollary 4.12 already established that $(\beta_{j})_{j\in S}$ is a consistent basis for $f|_{S}$. Therefore, by Lemma 4.15, the set $\beta_{a}$ contains a basis for $\mathrm{Im}[s,a]$ and $\mathrm{Ker}[a,s]$ for all $s\in S$. By hypothesis, every submodule of the form $\mathrm{Ker}[a,b]$ (respectively, $\mathrm{Im}[c,a]$) can be expressed in the form $\mathrm{Ker}[a,s]$ (respectively, $\mathrm{Im}[s,a]$) for some $s\in S$. The desired conclusion follows. ∎
@@
We relate the decomposability of integer persistent homology to the sensitivity of persistence diagrams to choice of coefficient field. Let $\mathcal{K}$ be a filtration, viewed as a functor from $I$ to the category of topological spaces and inclusions. Let $\mathcal{H}_{n}(\mathcal{K};F)$ be the functor obtained by composing $\mathcal{K}$ with the $n$th-homology functor over a field $F$. The persistence diagram of $\mathcal{K}$ with coefficients in $F$, denoted $\mathrm{PD}_{n}^{F}(\mathcal{K})$, is the multiset of intervals corresponding to an interval decomposition of $\mathcal{H}_{n}(\mathcal{K};F)$. The field-independence problem asks: under what conditions is the persistence diagram $\mathrm{PD}_{n}^{F}(\mathcal{K})$ independent of the choice of $F$? Obayashi and Yoshiwaki [12] addressed this in the special case where $I$ is finite. We extend this to the infinite case.
@@
(5.1) $\mathcal{H}_{n}(\mathcal{K}_{a};\mathbb{Z})$$\mathcal{H}_{n}(\mathcal{K}_{a};\mathbb{Z})\otimes F$$\mathcal{H}_{n}(\mathcal{K}_{a};F)$$\mathcal{H}_{n}(\mathcal{K}_{b};\mathbb{Z})$$\mathcal{H}_{n}(\mathcal{K}_{b};\mathbb{Z})\otimes F$$\mathcal{H}_{n}(\mathcal{K}_{b};{F})$$\scriptstyle{(*)}$$\scriptstyle{(\dagger)}\scriptstyle{\cong}$$\scriptstyle{(*)}$$\scriptstyle{(\dagger)}\scriptstyle{\cong}$
@@
    The module $\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})$ splits as a direct sum of interval submodules.

### v2

@@ 1. Introduction
Let $R$ be a principal ideal domain (PID). A persistence module with coefficients in $R$ is a functor $f:I\to R\mathrm{-Mod}$, where the indexing category $I$ is a totally-ordered poset category. Persistence modules [6, 9, 19] arise from and have been widely studied in persistent homology, in which setting $R$ is typically a field and $f$ maps every index $i\in I$ to a finite-dimensional vector space over $R$. In this setting, any persistence module admits an interval decomposition; that is, it can be decomposed as a direct sum of indecomposables known as “interval modules”. Interval decompositions are key to persistent homology, in that they capture important topological structure within data. If one removes the condition that the coefficient ring $R$ be a field, then interval decompositions are no longer guaranteed to exist [16, 7].
@@
The existence of interval decompositions for modules with field coefficients has been widely attributed to Gabriel [7], whose work showed that indecomposable representations of quivers of type $A_{n}$ have only finitely many isomorphism classes, which coincide exactly with interval modules. This result has since been generalized to persistence modules indexed by any totally-ordered set [4], as well as a small category [2]. [^1] The results in [4, 2] had been stated earlier (without proof) by Gabriel and Roĭter [8, Pages 30–32]. Interval decompositions have also been generalized to zigzag persistence. Carlson and de Silva [3] showed that interval decompositions exist for finitely-indexed zigzag persistence modules; Botnan [1] later showed the same in the infinitely-indexed setting. Igusa et al. [12] provided a proof for the more general continuous-zigzag setting, independent of [1]. Hang and Mio [11] generalized these results further, with interval decomposition of tame correspondence modules.
@@
Our work complements and builds on [13], which established necessary and sufficient conditions for interval decompositions of persistence modules over PIDs with finite indexing categories. The core element of [13] is a polynomial-time algorithm to compute decompositions explicitly. While key ideas in [13] remain relevant, the generality of our indexing category requires nontrivial adaptation. Roughly speaking, [13] constructs an interval decomposition of $f:\{1,\ldots,m\}\to R\mathrm{-Mod}$ by recursively constructing a basis for each $f_{i}$ using the basis on $f_{i-1}$. In the case were $I$ is infinite we would like to do something similar, but cannot proceed by exhaustion as in the finite case. Instead, we appeal to Zorn’s lemma, in a manner similar to [1]. As in that work, we lean heavily on the fact that, for fixed $i$, the set of distinct submodules of form $\mathrm{Ker}(f(i\leq j))$ (respectively, $\mathrm{Im}(f(k\leq k))$ has finite cardinality.
@@ 2.1. Persistence modules
We first show that the family $\{\mathrm{Ker}[a,y]\}_{y\in I}$ of kernels out of $f_{a}$ is finite. Let $y_{1},y_{2}\in I$ such that $a<y_{1}\leq y_{2}$. Then we have a short exact sequence
@@
We now show that the family $\{\mathrm{Im}[x,a]\}_{x\in I}$ of images into $f_{a}$ is finite. Let $x_{1},x_{2}\in I$ be such that $x_{1}\leq x_{2}<a$. Then we have a short exact sequence
@@ 2.2. A brief discussion on finitely-indexed persistence modules
If the structure maps of $f$ all have free cokernels, then [13, Theorem 17] states that, for each pair $(p,q)$ of integers such that $1\leq p,q\leq n$, there exits a (non-unique) submodule $A^{pq}_{a}\subseteq f_{a}$ such that
@@
Theorem [13, Theorem 18] states that, independent of one’s choice of $A^{pq}_{a}$, one has
@@
In the context of persistent homology, a submodule $A_{a}^{pq}\subseteq f_{a}$ in this decomposition can be thought of as cycles that are born at $p$ and die at $q$. This is because (1) whenever $a\notin[p,q)$, we must have $A_{a}^{pq}=0$ (see [13, Lemma 15]), and (2) one can choose submodules $A^{pq}_{a}$ such that $f(a\leq b)(A_{a}^{pq})=A_{b}^{pq}$. That is, for any choice of $A_{a}^{pq}$, the image $f(a\leq b)(A_{a}^{pq})$ is a complement of $\mathrm{IK}_{b}[p-1,q]+\mathrm{IK}_{b}[p,q-1]$ in $\mathrm{IK}_{b}[p,q]$ (see [13, Lemma 16]).
@@ 4. Sufficiency
Throughout this section, we fix a PID $R$, our coefficent ring; fix a persistence module $f:I\to R\mathrm{-Mod}$ that is nonzero (i.e., $f_{i}\neq 0$ for some $i$) and pointwise free and finitely-generated; assume that every structure map has free cokernel.
@@ 4.2. Partitioning consistent bases with images and kernels
Fix $J\subsetneq I$ and $a\in I\setminus J$. Fix a finite set $S_{J}$ such that $\{a\}\subseteq S_{J}\subseteq J\cup\{a\}$, which “covers” all images and kernels of $J$ in $f_{a}$, in the sense that $\{\mathrm{Im}[j,a]:j\in J\cup\{a\}\}=\{\mathrm{Im}[s,a]:s\in S_{J}\}$ and $\{\mathrm{Ker}[a,j]:j\in J\cup\{a\}\}=\{\mathrm{Ker}[a,s]:s\in S_{J}\}$. Such a finite set exists, by Lemma 2.9. Next, fix a finite set $S$ such that $S_{J}\subseteq S\subseteq I$, which covers all images and kernels of $I$ in $f_{a}$, in the sense that $\{\mathrm{Im}[i,a]:i\in I\}=\{\mathrm{Im}[s,a]:s\in S\}$ and $\{\mathrm{Ker}[a,i]:i\in I\}=\{\mathrm{Ker}[a,s]:s\in S\}$. Place the elements of $S$ into an ordered sequence $i_{1}<\cdots<i_{r}$.
@@ 4.3. Incremental Extension
    If none of the preceding rules applies (i.e., $i_{s}<i_{p}\leq a<i_{q}\leq i_{t}$), we choose any complement $A_{a}^{pq}$ of $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}[i_{p},i_{q-1}]$ in $\mathrm{IK}_{a}[i_{p},i_{q}]$, and define $\beta^{pq}_{a}$ to be any basis of $A_{a}^{pq}$. (Such a complement exists by [13, Theorem 17].) In this case, there is nothing in $J$ to push forward or pull back to index $a$, so we choose an arbitrary complement.
@@
In case (2), we have, by [13, Lemma 16], that $f(i_{s}\leq a)|_{{A_{i_{s}}^{pq}}}$ restricts to an isomorphism onto its image. Because $\beta^{pq}_{i_{s}}$ is linearly independent (it is a basis of $A^{pq}_{i_{s}}$), its image $f(i_{s}\leq a)(\beta^{pq}_{i_{s}})$ is as well.
@@
Case (1) Suppose that $i_{p}>a$ or $i_{q}\leq a$. By the proof of [13, Lemma 15], we have that $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]=\mathrm{IK}_{a}[i_{p},i_{q}]$. This implies that the complement is 0, with basis $\beta^{pq}_{a}:=\emptyset$, as desired.
Case (2) Suppose that $p\leq s$. By [13, Lemma 16], $f(i_{s}\leq a)(A_{i_{s}}^{pq})$ is a complement of $\mathrm{IK}_{a}[i_{p},i_{q-1}]+\mathrm{IK}_{a}[i_{p-1},i_{q}]$ in $\mathrm{IK}_{a}[i_{p},i_{q}]$.
@@
Figure 1. The four cases for defining $\beta_{a}^{pq}$, for different values of $i_{p}$ (birth) and $i_{q}$ (death). Red and blue segments show possible value ranges for $i_{q}$ and $i_{p}$, in each case. In Case (1) death occurs before birth, so $\beta_{a}^{pq}$ must be empty. In Case (2) birth occurs at or before $i_{s}$, so we may push some vectors from $\beta^{pq}_{i_{s}}$ forward to form $\beta^{pq}_{a}$. In Case (3) death occurs strictly after $i_{t}$, so we may pull some vectors back from $\beta^{pq}_{i_{t}}$. In Case (4) birth (respectively, death) occurs too late (respectively, too early), so no basis may be obtained from pushing forward or pulling back, and we must construct a new basis entirely.
@@
To show that $\beta_{a}=\bigsqcup_{p,q}\beta^{pq}_{a}$ is a basis for $f_{a}$, it is enough to show that $f_{a}=\bigoplus_{p,q}A^{pq}_{a}$. This is given by [13, Theorem 18]. ∎
@@
    $i_{p}>i_{s}$ or $i_{q}\leq i_{s}$. In this case, the basis $\beta^{pq}_{i_{s}}$ is empty, by [13, Lemma 15].
@@
Theorem 4.12 (Matching matrices for $(S\cap J)\cup\{a\}$).
@@
The matrix representation of every structure map with respect to the bases $(\beta_{j})_{j\in J\cup\{a\}}$ is a matching matrix. That is, for every $c,b\in J\cup\{a\}$ such that $c\leq b$, the matrix representation $f(c\leq b)]_{\beta_{c}}^{\beta_{b}}$ is a matching matrix.
@@
We first look at maps that end at $f_{a}$. Take any $c\in J$ such that $c<a$. Choose $i_{u}\in S\cap J$ such that $\mathrm{Im}[i_{u},a]=\mathrm{Im}[c,a]$. If $c\leq i_{u}\leq a$, then $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}=[f(i_{u}\leq a)]_{\beta_{i_{u}}}^{\beta_{a}}[f(c\leq i_{u})]_{\beta_{c}}^{\beta_{i_{u}}}$. Note that $f(c\leq i_{u})]_{\beta_{c}}^{\beta_{i_{u}}}$ is a matching matrix because $i_{u},c\in J$, and $(\beta_{j})_{j\in J}$ is a consistent basis. Additionally, by Theorem 4.12, $f(i_{u}\leq a)]_{\beta_{i_{u}}}^{\beta_{a}}$ is a matching matrix. Therefore $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix.
Now, consider $i_{u}\leq c<a$, and let $\{v_{1},\ldots,v_{r}\}:=\beta_{i_{u}}\setminus\mathrm{Ker}[i_{u},a]$. By Theorem 4.12, this set maps injectively into $\beta_{a}$, and its image under $f(i_{u}\leq a)$ spans $\mathrm{Im}[i_{u},a]=\mathrm{Im}[c,a]$. By functoriality, this implies that the image of the set $\{f(i_{u}\leq c)(v_{1}),\ldots,f(i_{u}\leq c)(v_{r})\}\subseteq\beta_{c}$ maps injectively into the basis $\beta_{a}\subseteq f_{a}$ under $f(c\leq a)$. Because $\beta_{c}$ contains a basis for $\mathrm{Ker}[c,a]$, we have, by the rank-nullity theorem, that $\beta_{c}\setminus\{f(i_{u}\leq c)(v_{1}),\ldots,f(i_{u}\leq c)(v_{r})\}$ spans $\mathrm{Ker}[c,a]$. Therefore, it follows that $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix.
Next, we look at maps that begin at $f_{a}$. Take any $b\in J$ such that $b>a$. By definition, there exists $i_{\ell}\in S\cap J$ such that $\mathrm{Ker}[a,i_{\ell}]=\mathrm{Ker}[a,b]$. If $a<i_{\ell}\leq b$, then $f(a\leq b)]_{\beta_{a}}^{\beta_{b}}=[f(i_{\ell}\leq b)]_{\beta_{i_{\ell}}}^{\beta_{b}}[f(a\leq i_{\ell})]_{\beta_{a}}^{\beta_{i_{\ell}}}$ is a matching matrix because both $f(i_{\ell}\leq b)]_{\beta_{i_{\ell}}}^{\beta_{b}}$ and $f(a\leq i_{\ell})]_{\beta_{a}}^{\beta_{i_{\ell}}}$ (by Theorem 4.12) are matching matrices.
Now, suppose that $a<b\leq i_{\ell}$. Let $v\in\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]$ be given. Because $\mathrm{Ker}[a,i_{\ell}]=\mathrm{Ker}[a,b]$, the elements $f(a\leq b)(v)$ and $f(a\leq i_{\ell})(v)$} are nonzero. We know that $f(a\leq i_{\ell})(v)\in\beta_{i_{\ell}}$, by Theorem 4.12. Therefore, because $(\beta_{j})_{j\in J}$ is a consistent basis, there exists $w\in\beta_{b}$ such that $f(b\leq i_{\ell})(w)=f(a\leq i_{\ell})(v)$.
@@
From this, it follows that $f(a\leq b)$ maps $\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]=\beta_{a}\setminus\mathrm{Ker}[a,b]$ into $\beta_{b}$. Moreover, functoriality implies that this map is injective, because the composite map $f(a\leq i_{\ell})=f(b\leq i_{\ell})\circ f(a\leq b)$ is injective on $\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]=\beta_{a}\setminus\mathrm{Ker}[a,b]$, by Theorem 4.12. This implies that $f(a\leq b)]_{\beta_{a}}^{\beta_{b}}$ is a matching matrix, as desired. ∎
@@
Suppose that $f:I\to R\mathrm{-Mod}$ is any persistence module with consistent basis $(\beta_{i})_{i\in I}$. Then, for any $a\in I$, the basis $\beta_{a}$ contains a basis for every submodule of the form $\mathrm{Im}[c,a]$ and $\mathrm{Ker}[a,b]$.
@@
For any $x,y\in I$ such that $x\leq y$, define $\mathcal{K}_{x,y}=\beta_{x}\cap\mathrm{Ker}[x,y]$, and $\mathcal{I}_{x,y}=\beta_{x}\setminus\mathcal{K}_{x,y}$. Let $c\leq a\leq b$ be given. Clearly, $f(c\leq a)(\mathcal{I}_{c,a})$ spans the image submodule $\mathrm{Im}[c,a]$. Because $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix, the set $\mathcal{I}_{c,a}$ maps injectively into $\beta_{a}$, and therefore, $f(c\leq a)(\mathcal{I}_{c,a})\subseteq\beta_{a}$ forms a basis for $\mathrm{Im}[c,a]$. Moreover, by the rank nullity theorem, $\mathcal{K}_{a,b}$ forms a basis for $\mathrm{Ker}[a,b]$. ∎
@@
Theorem 4.12 already established that $(\beta_{j})_{j\in S}$ is a consistent basis for $f|_{S}$. Therefore, by Lemma 4.15, the set $\beta_{a}$ contains a basis for $\mathrm{Im}[s,a]$ and $\mathrm{Ker}[a,s]$ for all $s\in S$. By hypothesis, every submodule of the form $\mathrm{Ker}[a,b]$ (respectively, $\mathrm{Im}[c,a]$) can be expressed in the form $\mathrm{Ker}[a,s]$ (respectively, $\mathrm{Im}[s,a]$) for some $s\in S$. The desired conclusion follows. ∎
@@
We relate the decomposability of integer persistent homology to the sensitivity of persistence diagrams to choice of coefficient field. Let $\mathcal{K}$ be a filtration, viewed as a functor from $I$ to the category of topological spaces and inclusions. Let ${H}_{n}(\mathcal{K};F)$ be the functor obtained by composing $\mathcal{K}$ with the $n$th-homology functor over a field $F$. The persistence diagram of $\mathcal{K}$ with coefficients in $F$, denoted $\mathrm{PD}_{n}^{F}(\mathcal{K})$, is the multiset of intervals corresponding to an interval decomposition of ${H}_{n}(\mathcal{K};F)$. The field-independence problem asks: under what conditions is the persistence diagram $\mathrm{PD}_{n}^{F}(\mathcal{K})$ independent of the choice of $F$? Obayashi and Yoshiwaki [15, Theorem 1.6] proved that, when $I$ is finite, a persistence diagram of a filtration is independent of field choice if and only if the relative homology group $H_{n}(\mathcal{K}_{p},\mathcal{K}_{q};\mathbb{Z})$ is free for all $q\leq p$, and $H_{n-1}(\mathcal{K}_{p};\mathbb{Z})$ is free for all $p$. We extend this to the infinite case.
@@
(5.1) $H_{n}(\mathcal{K}_{a};\mathbb{Z})$$H_{n}(\mathcal{K}_{a};\mathbb{Z})\otimes F$$H_{n}(\mathcal{K}_{a};F)$$H_{n}(\mathcal{K}_{b};\mathbb{Z})$$H_{n}(\mathcal{K}_{b};\mathbb{Z})\otimes F$$H_{n}(\mathcal{K}_{b};{F})$$\scriptstyle{(*)}$$\scriptstyle{(\dagger)}\scriptstyle{\cong}$$\scriptstyle{(*)}$$\scriptstyle{(\dagger)}\scriptstyle{\cong}$
@@
    The module $H_{n}(\mathcal{K};\mathbb{Z})$ splits as a direct sum of interval submodules.